# Dust Properties of the Interstellar Object 3I/ATLAS Revealed by Optical and Near-Infrared Polarimetry

### v1

@@ Abstract
We present independent polarimetric observations of the interstellar object 3I/ATLAS, including the first near-infrared polarimetric measurements. Using imaging polarimeters, we measured the degree of linear polarization from the visible $R_{\mathrm{C}}$band (0.64 µm) to the near-infrared $K_{\mathrm{s}}$band (2.25 µm), and investigated its dependence on solar phase angle (polarization phase curve; PPC) and wavelength (polarization color curve; PCC). We confirm that the PPC of 3I/ATLAS differs significantly from those of typical Solar System comets, showing an unusually large polarization amplitude. This PPC shows no significant change in the $R_{C}$ band across perihelion passage, despite the perihelion lying within the water snow line. This indicates that the unusual polarimetric behavior of 3I/ATLAS is unlikely to be driven by transient volatile activity, but instead reflects intrinsic optical properties of refractory dust particles. The PCC increases with wavelength over 0.6–1.2 µm and peaks at 1.5–2.0 µm, suggesting that the dominant scattering units are dust aggregates composed of submicron-sized monomers, broadly consistent with interstellar dust and solar-system cometary aggregates. Taken together, our results indicate that 3I/ATLAS preserves polarimetric properties characteristic of a primitive cometary planetesimal formed in another planetary system, with a refractory dust composition that differs from that typically observed among Solar System comets, despite sharing a similar size scale of the aggregate building blocks.
@@ I Introduction
The recent discovery of ISOs offers a unique opportunity to test whether this apparent uniformity of cometary dust properties is a universal outcome of planet formation or a peculiarity of the Solar System. Polarimetric observations of the second ISO, 2I/Borisov, revealed subtle but measurable deviations from the typical SSC polarization phase curves. On 2025 July 1, the third known ISO, 3I/ATLAS, was discovered at a heliocentric distance of 4.49 au. Soon after its discovery(Denneau et al., 2025), intensive polarimetric observations were initiated by Gray et al. (2025). They reported that the polarization degree of 3I/ATLAS differs significantly from that of SSCs. Furthermore, Zheltobryukhov et al. (2025) reported an apparent discontinuous change in its PPC across the water snow line, and suggested that the presence of volatile-rich material may play a key role in shaping the unusual polarimetric behavior of this object.
@@ III Results
Figure 1 shows the PPCs of 3I/ATLAS measured in the $R_{\mathrm{C}}$, $I_{\mathrm{C}}$, $J$-, $H$- and $K_{\mathrm{s}}$-bands. The $R_{\mathrm{C}}$-band observations provide the highest sampling frequency, covering a phase-angle range of approximately $\alpha\simeq 19\arcdeg$$32\arcdeg$ during post-perihelion orbital phase. For comparison, previously published polarimetric data of 3I/ATLAS (Gray et al. 2025, obtained at heliocentric distances of $r_{H}\simeq 2.6$–4.0 au, and Zheltobryukhov et al. 2025, obtained at $r_{H}\simeq 1.4$–1.7 au), as well as PPCs of other SSCs, are also shown in the same figure. The latter observations were conducted at heliocentric distances where the ambient temperature is sufficiently higher than the sublimation temperature of water ice, corresponding qualitatively to the region interior to the water snow line.
@@
The $I_{\mathrm{C}}$-band PPC shows behavior broadly consistent with that in the $R_{\mathrm{C}}$-band. At near-infrared wavelengths ($J$, $H$, and $K_{\mathrm{s}}$bands), the degree of polarization increases monotonically with phase angle over the observed range. However, due to the limited number of measurements at small phase angles ($\alpha\lesssim 23\arcdeg$), it is difficult to determine an inversion angle at these wavelengths. As in the $R_{\mathrm{C}}$-band case, the depth of the NPB at all observed wavelengths is significantly deeper than that of typical SSCs, indicating that the polarimetric behavior of 3I/ATLAS is distinctive across all observed wavelengths.
Furthermore, the PPCs obtained during the inbound and outbound phases are mutually consistent within the observational uncertainties in all bands, and no significant temporal variation in the polarimetric properties is detected over the observed period. These measurements were obtained at comparable heliocentric distances for a given phase angle, allowing a direct comparison between the two orbital phases.
@@
Figure 2: Spectral dependence of the polarization degree of 3I/ATLAS at $\alpha=23.4\arcdeg$ and $30.5\arcdeg$, compared with those of comet C/Hale–Bopp, C/Tsuchinshan–ATLAS, and 1P/Halley. For the comparison comets, polarization degrees in each band are taken from the literature at similar phase angles and, when necessary, are estimated by extrapolation (Appendix B; Kikuchi et al. 1987; Kiselev et al. 2017; Lim et al. 2025). The dashed and solid curves show Chebyshev polynomial fits.
@@ IV.4 Implications for the Nature of Dust in 3I/ATLAS
This research was supported by a grant from the Korean National Research Foundation (NRF) (MEST), funded by the Korean government (No. 2023R1A2C1006180).
@@ Appendix A Appendix: Data Reduction and Error Analysis
UT Date $\alpha$[°] Instrument Filter $P$ [%] $\sigma_{P}$ [%] $\theta$ [°] $\sigma_{\theta}$ [°] $P_{\mathrm{r}}$ [%] $\theta_{\mathrm{r}}$ [°]
@@
2025-12-03.86 30.54 SQUIDPOL $I_{\mathrm{C}}$ 6.93 0.24 24.61 0.99 6.92 0.98
2025-12-08.83 29.37 SQUIDPOL $R_{\mathrm{C}}$ 6.71 0.16 19.65 0.70 6.65 -3.60
@@
2025-12-21.80 21.95 SQUIDPOL $R_{\mathrm{C}}$ 2.60 0.18 22.98 2.03 2.59 1.01
2025-12-25.81 19.70 SQUIDPOL $R_{\mathrm{C}}$ 1.34 0.55 20.41 11.75 1.23 -1.19
2026-01-07.80 9.66 HONIR $R_{\mathrm{C}}$ 2.21 0.17 -12.94 2.25 -2.19 86.43
@@
2026-01-08.54 9.10 MSI $R_{\mathrm{C}}$ 2.31 0.23 111.83 2.90 -2.30 91.13
The preprocessing includes bias subtraction, dark subtraction, and flat-field correction, together with cosmic-ray rejection. For NIC data, the vertical pattern noise was corrected at the raw-image level using the NIC data reduction package NICpolpy (Bach et al., 2022). After preprocessing, fluxes were extracted using aperture photometry. In the case of SQUIDPOL, the large seeing and tracking errors broadened the point-spread function (PSF) of 3I/ATLAS. Therefore, we adopted an aperture radius of approximately 30 pixels to collect as much signal from 3I/ATLAS as possible, which physically corresponds to about $1.5\times 10^{4}-2.0\times 10^{4}$km and thus includes light not only from the nucleus but also from the surrounding coma.
@@ Appendix B PCC Estimation for C/Hale–Bopp and C/Tsuchinshan–ATLAS
In Figure 2, the data points for two comets, C/Hale–Bopp and C/Tsuchinshan–ATLAS, are extrapolated values, because no polarimetric measurements are available at phase angles of $\alpha\sim 23\arcdeg$ and $\alpha\sim 30\arcdeg$ for these comets. Using Equation (1), we constructed polarization phase curves (PPCs) in each filter ($R_{\mathrm{C}}$, $I_{\mathrm{C}}$, $J$, $H$, and $K_{\mathrm{s}}$) and extrapolated the corresponding polarization degrees $P_{\mathrm{r}}$ at the required phase angles.
Since Equation (1) reliably reproduces observed polarization behavior only at small phase angles, the fitting was performed using data points with phase angles smaller than $70\arcdeg$. Figures 3 and 4 show the resulting PPC fits, where the red and blue symbols indicate the extrapolated polarization degrees at $\alpha=23.4\arcdeg$ and $\alpha=30.5\arcdeg$, respectively.
Figure 3: PPC of C/Hale-Bopp within each filter
Figure 4: PPC of C/Tsuchinshan-ATLAS within each filter

### v2

@@ Abstract
We present independent polarimetric observations of the interstellar object 3I/ATLAS, including the first near-infrared polarimetric measurements. Using imaging polarimeters, we measured the degree of linear polarization from the visible $R_{\mathrm{C}}$-band (0.64 µm) to the near-infrared $K_{\mathrm{s}}$-band (2.25 µm), and investigated its dependence on solar phase angle (polarization phase curve; PPC) and wavelength (polarization color curve; PCC). We confirm that the PPC of 3I/ATLAS differs significantly from those of typical Solar System comets, showing an unusually large polarization amplitude. This PPC shows no significant change in the $R_{\mathrm{C}}$ band across perihelion passage, despite the perihelion lying within the water snow line. This indicates that the unusual polarimetric behavior of 3I/ATLAS is unlikely to be driven by transient volatile activity, but instead reflects intrinsic optical properties of refractory dust particles. The PCC increases with wavelength over 0.6–1.2 µm and peaks at 1.5–2.0 µm, suggesting that the dominant scattering units are dust aggregates composed of submicron-sized monomers, broadly consistent with interstellar dust and solar-system cometary aggregates. Taken together, our results indicate that 3I/ATLAS preserves polarimetric properties characteristic of a primitive cometary planetesimal formed in another planetary system, with a refractory dust composition that differs from that typically observed among Solar System comets, despite sharing a similar size scale of the aggregate building blocks.
@@ I Introduction
The recent discovery of ISOs offers a unique opportunity to test whether this apparent uniformity of cometary dust properties is a universal outcome of planet formation or a peculiarity of the Solar System. Polarimetric observations of the second ISO, 2I/Borisov, revealed subtle but measurable deviations from the typical SSC polarization phase curves (Bagnulo et al., 2021). On 2025 July 1, the third known ISO, 3I/ATLAS, was discovered at a heliocentric distance of 4.49 au. Soon after its discovery (Denneau et al., 2025), intensive polarimetric observations were initiated by Gray et al. (2025). They reported that the polarization degree of 3I/ATLAS differs significantly from that of SSCs before crossing the water snow line. Furthermore, Zheltobryukhov et al. (2025) reported an apparent discontinuous change in its PPC across the water snow line, and suggested that the presence of volatile-rich material may play a key role in shaping the unusual polarimetric behavior of this object.
@@ III Results
Figure 1 shows the PPCs of 3I/ATLAS measured in the $R_{\mathrm{C}}$, $I_{\mathrm{C}}$, $J$-, $H$- and $K_{\mathrm{s}}$-bands. The $R_{\mathrm{C}}$-band observations provide the highest sampling frequency, covering a phase-angle range of approximately $\alpha\simeq 9\arcdeg$$32\arcdeg$ during post-perihelion orbital phase. For comparison, previously published polarimetric data of 3I/ATLAS (Gray et al. 2025, obtained at heliocentric distances of $r_{H}\simeq 2.6$–4.0 au before crossing the water snow line, and Zheltobryukhov et al. 2025, obtained at $r_{H}\simeq 1.4$–1.7 au after crossing the water snow line), as well as PPCs of other SSCs, are also shown in the same figure. The latter observations were conducted at heliocentric distances where the ambient temperature is sufficiently higher than the sublimation temperature of water ice, corresponding qualitatively to the region interior to the water snow line.
@@
The $I_{\mathrm{C}}$-band PPC shows behavior broadly consistent with that in the $R_{\mathrm{C}}$-band. At near-infrared wavelengths ($J$, $H$, and $K_{\mathrm{s}}$-bands), the degree of polarization increases monotonically with phase angle over the observed range. However, due to the limited number of measurements at small phase angles ($\alpha\lesssim 23\arcdeg$), it is difficult to determine an inversion angle at these wavelengths. As in the $R_{\mathrm{C}}$-band case, the depth of the NPB at all observed wavelengths is significantly deeper than that of typical SSCs, indicating that the polarimetric behavior of 3I/ATLAS is distinctive across all observed wavelengths.
Furthermore, the PPCs obtained during the inbound from (Gray et al., 2025) and outbound phases from this study are mutually consistent within the observational uncertainties in all bands, and no significant temporal variation in the polarimetric properties is detected over the observed period. These measurements were obtained at comparable heliocentric distances for a given phase angle, allowing a direct comparison between the two orbital phases.
@@
Figure 2: Spectral dependence of the polarization degree of 3I/ATLAS at $\alpha=24.2\arcdeg$$24.9\arcdeg$ and $30.5\arcdeg$, compared with those of comet C/Hale–Bopp, C/Tsuchinshan–ATLAS, and 1P/Halley. For the comparison comets, polarization degrees in each band are taken from the literature at similar phase angles and, when necessary, are estimated by extrapolation (Appendix B; Kikuchi et al. 1987; Kiselev et al. 2017; Lim et al. 2025). The dashed and solid curves show Chebyshev polynomial fits.
@@ IV.1 Reliability of our data
In addition, we considered the possible influence of gas depolarization in our data. Based on the spectroscopic measurements by Hoogendam et al. (2026), the CN emission contributes approximately 1% of the total $R_{\mathrm{C}}$-band flux around November 16. This small line-to-continuum ratio is broadly consistent with the spectrum obtained on 2025 August 15 by SPHEREx (Lisse et al., 2025), although the observing epochs are slightly different. Although CN emission may possess some intrinsic polarization (e.g., Kwon et al., 2018), it is generally much lower than that of dust. Assuming that the CN emission is completely unpolarized, we estimate a conservative upper limit on the depolarization effect. In that case, the maximum depolarization would be only $\sim$0.04% (i.e., $0.01\times 4.32\%$), which is significantly smaller than our typical measurement uncertainty of $\sim$0.2%. Although other emission species (e.g., NH2) may also contribute in the $R_{\mathrm{C}}$-band, CN is generally brighter than NH2 (Kwon et al., 2018), suggesting that their contribution is even smaller. Therefore, gas contamination in the $R_{\mathrm{C}}$ band can safely be considered negligible.
For the $I_{\mathrm{C}}$band, the red CN system may be relatively stronger at longer wavelengths (Kwon et al., 2018). While a minor depolarization effect cannot be entirely excluded in the $I_{\mathrm{C}}$band, the polarization degrees in this band appear slightly lower than the expected phase-angle trend near perihelion (Figure 1). Nevertheless, for most of our observations, the overall gas contribution is expected to remain small compared to the measured polarization amplitude.
@@ IV.4 Implications for the Nature of Dust in 3I/ATLAS
V Summary
We present the post-perihelion polarimetric data of 3I/ATLAS, covering a phase angle range of $\alpha\simeq 9\arcdeg$$32\arcdeg$ in both optical ($R_{\mathrm{C}}$ and $I_{\mathrm{C}}$) and near-infrared ($J$-, $H$-, and $K_{\mathrm{s}}$-bands) wavelengths. Our optical polarimetric results were cross-calibrated between the SQUIDPOL instrument and the Pirka/MSI system. The key findings are as follows: (i) the PPC exhibits a large polarization amplitude and shows good consistency with the pre-perihelion data (Gray et al., 2025), and (ii) the PCC displays behavior similar to Solar System comets. These results suggest that 3I/ATLAS consists of refractory grains with an aggregate structure, similar in size to typical Solar System comets. Combining the PPC and PCC results, we interpret these findings as indicating that, although compositional differences may exist in refractory dust, the fundamental size of the scattering units may be broadly similar to that inferred for typical Solar System comets.
Data Availability
The availability of the observational data listed in Table 1 is as follows. (1) SQUIDPOL data: The polarimetric data products are publicly available via Zenodo [^1] https://zenodo.org/records/18627153 with a DOI (10.5281/zenodo.18627153). The archive includes the calibration data, observational data of 3I/ATLAS, and associated metadata necessary for reuse. (2) Pirka Telescope data: The Pirka observations are not stored in a public archive. The observational data products are available upon reasonable request to the co-authors: Hisayuki Kubota (hkubota@sci.hokudai.ac.jp) and Seiko Takagi (seiko@ep.sci.hokudai.ac.jp). (3) Nishiharima Astronomical Observatory and Higashi-Hiroshima Observatory data: These observational data will be made publicly available through the SMOKA (Subaru–Mitaka–Okayama–Kiso Archive System) archive after the standard 18-month proprietary period. Readers are encouraged to access the data via SMOKA once they are released (Baba et al., 2002).
This research was supported by a grant from the Korean National Research Foundation (NRF) (MEST), funded by the Korean government (No. 2023R1A2C1006180). We thank the anonymous referee for constructive comments that helped improve the clarity of the manuscript.
@@ Appendix A Appendix: Data Reduction and Error Analysis
UT Date $\alpha$[°] $r_{H}$[au] Instrument Filter $P$ [%] $\sigma_{P}$ [%] $\theta$ [°] $\sigma_{\theta}$ [°] $P_{\mathrm{r}}$ [%] $\theta_{\mathrm{r}}$ [°]
@@
2025-12-03.86 30.44 1.87 SQUIDPOL $I_{\mathrm{C}}$ 6.93 0.24 24.61 0.99 6.92 0.98
2025-12-08.83 29.37 2.00 SQUIDPOL $R_{\mathrm{C}}$ 6.71 0.16 19.65 0.70 6.65 -3.60
@@
2025-12-21.80 22.68 2.36 SQUIDPOL $R_{\mathrm{C}}$ 2.60 0.18 22.98 2.03 2.59 1.01
2025-12-25.81 19.70 2.48 SQUIDPOL $R_{\mathrm{C}}$ 1.34 0.55 20.41 11.75 1.23 -1.19
2026-01-07.80 9.66 2.88 HONIR $R_{\mathrm{C}}$ 2.21 0.17 -12.94 2.25 -2.19 86.43
@@
2026-01-08.54 9.10 2.90 MSI $R_{\mathrm{C}}$ 2.31 0.23 111.83 2.90 -2.30 91.13
The preprocessing includes bias subtraction, dark subtraction, and flat-field correction, together with cosmic-ray rejection. For NIC data, the vertical pattern noise was corrected at the raw-image level using the NIC data reduction package NICpolpy (Bach et al., 2022). After preprocessing, fluxes were extracted using aperture photometry. In the case of SQUIDPOL, the large seeing and tracking errors broadened the point-spread function (PSF) of 3I/ATLAS. Therefore, we adopted an aperture radius of approximately 30 pixels to collect as much signal from 3I/ATLAS as possible, which physically corresponds to about $1.5\times 10^{4}-2.0\times 10^{4}$ km at the location of the comet and thus includes light not only from the nucleus but also from the surrounding coma. This aperture size was confirmed with the analysis based on aperture size dependency on Figure A1.
Figure A1: Aperture size dependency of polarimetric analysis for observed data on Nov.14
@@ Appendix B PCC Estimation for C/Hale–Bopp and C/Tsuchinshan–ATLAS
In Figure 2, the data points for two comets, C/Hale–Bopp and C/Tsuchinshan–ATLAS, are extrapolated values, because no polarimetric measurements are available at phase angles of $\alpha\sim 24\arcdeg$ and $\alpha\sim 30\arcdeg$ for these comets. Using Equation (1), we constructed polarization phase curves (PPCs) in each filter ($R_{\mathrm{C}}$, $I_{\mathrm{C}}$, $J$, $H$, and $K_{\mathrm{s}}$) and extrapolated the corresponding polarization degrees $P_{\mathrm{r}}$ at the required phase angles.
Since Equation (1) reliably reproduces observed polarization behavior only at small phase angles, the fitting was performed using data points with phase angles smaller than $70\arcdeg$. Figures B1 and B2 show the resulting PPC fits, where the red and blue symbols indicate the extrapolated polarization degrees at $\alpha=24.5\arcdeg$ and $\alpha=30.5\arcdeg$, respectively.
Figure B1: PPC of C/Hale-Bopp within each filter
Figure B2: PPC of C/Tsuchinshan-ATLAS within each filter